OpenStudy (anonymous):

f(t)=sin t is an odd function g(t)=cos t even function what can be said about h(t)=f(t)g(t) ??????

OpenStudy (anonymous):

h(-t) = f(-t) g(-t) = -f(t) g(t) = -h(t) so h(t) is odd

OpenStudy (anonymous):

omg ily THANKS

OpenStudy (anonymous):

this is true no matter what f and g are. notice that sin(t) cos(t) = (1/2) sin(2t), which is odd

OpenStudy (anonymous):

you're like the only one that is like replying i've been answering my self for the past 2 questions fjdfhksafhkashfkjsd

OpenStudy (anonymous):

KAY THANKS